Some recent financial statements for Smolira Golf, Inc., follow.

SMOLIRA GOLF, INC.
Balance Sheets as of December 31, 2015 and 2016
2015 2016 2015 2016
Assets Liabilities and Owners’ Equity
Current assets Current liabilities
Cash $ 3,121 $ 3,157 Accounts payable $ 2,168 $ 2,630
Accounts receivable 4,752 5,751 Notes payable 1,765 2,146
Inventory 12,538 13,752 Other 93 110
Total $ 20,411 $ 22,660 Total $ 4,026 $ 4,886
Long-term debt $ 13,100 $ 15,860
Owners’ equity
Common stock and paid-in surplus $ 39,500 $ 39,500
Fixed assets Accumulated retained earnings 15,669 39,331
Net plant and equipment $ 51,884 $ 76,917 Total $ 55,169 $ 78,831
Total assets $ 72,295 $ 99,577 Total liabilities and owners’ equity $ 72,295 $ 99,577
SMOLIRA GOLF, INC.
2016 Income Statement
Sales $ 187,970
Cost of goods sold 126,503
Depreciation 5,303
EBIT $ 56,164
Interest paid 1,400
Taxable income $ 54,764
Taxes 19,167
Net income $ 35,597
Dividends $ 11,935
Retained earnings 23,662

Find the following financial ratios for Smolira Golf (use year-end figures rather than average values where appropriate): (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16. Enter the profitability ratios as a percent.)

2015 2016
Short-term solvency ratios
a. Current ratio times times
b. Quick ratio times times
c. Cash ratio times times
Asset utilization ratios
d. Total asset turnover times
e. Inventory turnover times
f. Receivables turnover times
Long-term solvency ratios
g. Total debt ratio times times
h. Debt−equity ratio times times
i. Equity multiplier times times
j. Times interest earned ratio times
k. Cash coverage ratio times
Profitability ratios
l. Profit margin %
m. Return on assets %
n. Return on equity %

PLEASE DO F-N
Answer #1

For 2016:

Receivables Turnover ratio = net credit sales/ average accounts receivable

net credit sales = 187,970

average accounts receivable = (opening accounts receivable + closing accounts receivable)/ 2

= (4752+5751)/2 = 5251.5

Receivables Turnover ratio = 187,970/5251.5 = 35.79

Total debt ratio = total debt/ total assets

Total debt ratio for 2015 = 13,100/ 72,295 = 0.1812 = 18.12%

Total debt ratio for 2016 = 15,860/ 99,577 = 0.1592 = 15.92%

debt equity ratio = total debt/ total equity

debt equity ratio for 2015 = 13,100/ 55,169 = 0.23745 = 23.75%

debt equity ratio for 2016 = 15,860/ 78,831 = 0.20118 = 20.12%

equity multiplier = total assets/ total shareholders equity

equity multiplier for 2015 = 72,295/ 55,169 = 1.31

equity multiplier for 2016 = 99,577/ 78,831 = 1.26

For 2016:

Time interest earned ratio = EBIT/ interest paid = 56,164/ 1,400 = 40.12

cash coverage ratio = (cash + cash equivalents)/ total current liabilities = (3,157 + 5,751)/ 4,886 = 1.82

profit margin = net income/ net sales = 35,597/ 187,970 = 0.18937 = 18.94%

return on assets = net income/ average total assets

average total assets = (72,295 + 99,577)/ 2 = 85,936

return on assets = 35,597/ 85,936 = 0.41422 = 41.42%

return on equity = net income/ average equity

average equity = (55,169 + 78,831)/ 2 = 67,000

return on equity = 35,597/ 67,000 = 0.53129 = 53.13%